OpenStudy (anonymous):

What is the repeating decimal patternfor fractions with a denominator of 13?

OpenStudy (anonymous):

what is \(\frac{1}{13}\)?

OpenStudy (anonymous):

0.076923076923.......

OpenStudy (unklerhaukus):

\[x=\frac 1{13}=0.07692307692307693....\]\[\quad=0.0\dot769230\dot7\] \[1000000x=76923.0\dot769230\dot7\] \[999999x=76923\] \[x=\frac{76923}{999999}\] \[\frac 1{13}=\frac{76923}{999999}\] \[\frac n{13}=n\frac{76923}{999999}\]

OpenStudy (mayankdevnani):

repeating decimal is that___ 1/13=0.076923076923076923

OpenStudy (mayankdevnani):

that is 0.076923

OpenStudy (mayankdevnani):

@yociyoci

OpenStudy (anonymous):

but i am trying to find the decimal pattern for any fractions with the denominator of 13, like 1/13 = 0.076923076923 and 2/13 = 0.153846153846 and 3/13= 0.230769230769 by analyzing these, how do you find the pattern for the next ones?

OpenStudy (mayankdevnani):

by multiplying 0.076923 by number what you want like 9/13=? then you would multiply 0.076923 by 9 to get =9/13

OpenStudy (mayankdevnani):

@yociyoci

OpenStudy (mayankdevnani):

9/13 = 0.692307 whole bar... on 692307 not on 0

OpenStudy (anonymous):

so there isn't a number pattern to the decimal from 1/13 to 12/13?

OpenStudy (unklerhaukus):

there is a pattern

OpenStudy (anonymous):

Can you please clarify?

OpenStudy (mayankdevnani):

yaaa!! there is a pattern that multiply 0.076923 by 12 to get =12/13

OpenStudy (anonymous):

ok, sure

OpenStudy (anonymous):

thx

OpenStudy (mayankdevnani):

welcome...